Complete this table of values for three aqueous solutions at 25 °C он pH Given Number Number Solution A: Number Given Number Solution B: 0.0045 Number Number Given Solution C: | | | 7.49
Answer #1

ANSWER:

we know that,

pH = -log[H+]

[H+][OH-] = 10-14

For solution A:

[H+] = 0.00055 M

pH = - log[H+]

pH = -log(0.00055)

pH= 3.26

[H+][OH-] = 10-14

(0.00055 M)[OH-] = 10-14

[OH-] = 1.82 x 10-11 M

For solution B:

[OH-] = 0.0045 M

[H+][OH-] = 10-14

[H+](0.0045 M) = 10-14

[H+] = 2.22 x 10-12 M

pH = -log[H+]

pH = -log(2.22 x 10-12 M)

pH = 11.65

For solution C:

pH = 7.49

pH = -log[H+]

7.49 = -log[H+]

[H+] = 3.24 x 10-8 M

[H+][OH-] = 10-14

(3.24 x 10-8 M)[OH-] = 10-14

[OH-] = 3.09 x 10-7 M

[H+] [OH-] pH
Solution A 0.00055 M 1.82 x 10-11 M 3.26
Solution B 2.22 x 10-12 M 0.0045 M 11.65
Solution C 3.24 x 10-8 M 3.09 x 10-7 M 7.49